# Pooling Liquidity Pools in AMMs

### v1

@@ 4 The Analysis of Global Market Makers
for $x_{i}\equiv x_{i}^{0}+\Delta x_{i}$ and $y_{i}\equiv y_{i}^{0}-\Delta y_{i}$. Now suppose that there is a sequence of swaps submitted to the other AMMs that leads to a new vector of reserves $(x_{-i},y_{-i})$ such that:
$\frac{y_{-i}}{x_{-i}}\leq\frac{y_{i}}{x_{i}},$ (7)
and consider a swap that sends $\Delta y_{i}$ of $Y$ to AMM $i$. Then:
$\Delta x_{CPMM}(\Delta y_{i};x_{i},y_{i})$ $=$ $\frac{\frac{x_{i}}{y_{i}}}{1+\frac{\Delta y_{i}}{y_{i}}}\Delta y_{i}$
$\leq$ $\frac{\frac{x}{y}}{1+\frac{\Delta y_{i}}{y}}\Delta y_{i}$
$=$ $\Delta x_{nGMM}(\Delta y_{i};x,y),$
where the two equalitites follow from (2) and (4), and the inequality from the fact that $x_{i}<x$ and that (7) implies that $\frac{y_{i}}{x_{i}}\geq\frac{y}{x}$. This together with our restriction to global algorithms between the CPMM and the nGMM means that the amount of $X$ that AMM $i$ gives to the trader is at least:
$\hat{\Delta}x_{i}\geq\Delta x_{CPMM}(\Delta x_{i};x_{i},y_{i}),$ (8)
which by the definition of $\Delta x_{CPMM}(\Delta x_{i};x_{i},y_{i})$ implies that:
$(x_{i}-\hat{\Delta}x_{i})(y_{i}+\Delta y)\leq x_{i}y_{i}.$
Since $y_{i}=y^{0}_{i}-\Delta y_{i}$, (6) means that $x_{i}-\hat{\Delta}x_{i}<x_{i}^{0}$ and thus the updated vector of reserves of AMM $i$ satisfies,
$(x_{i}-\hat{\Delta}x_{i},y^{0}_{i})\leq(x_{i}^{0},y^{0}_{i}),$
as desired. ∎
Thus, global algorithms in between the CPMM and nGMM buy asset $X$ at a price no larger than the CPMM price to avoid exploitation. Intuitively, the CPMM price guaranties that when the AMM buys an asset and sells it back the vector of reserves returns to its initial level. Thus, whenever a global algorithm buys one of the assets more expensive than the CPMM it is going to be exploited if it sells it back at CPMM prices or lower. But this will always happen with global algorithms with prices in between the CPMM and nGMM if the asset becomes so abundant among the other AMMs between the initial sale and the sell back that the nGMM price becomes lower than the CPMM price.
Furthermore, since the CPMM is not exhaustible and Proposition 2 says that to avoid exploitation the price at which the algorithm buys must be at most the CPMM price, we can conclude that the algorithms that satisfy Proposition 2 are not exhaustible.
Corollary 1.
A non exploitable global algorithm between the CPMM and the nGMM is also non exhaustible.
Since CPMM prices put a lower bound to global non exploitable algorithms, the closer one can get to nGMM prices while avoiding exploitation (and thus exhaustability) is to use nGMM prices unless bound by the CPMM.
@@
To understand the GMM, it is useful to distinguish whether a swap $\Delta x_{i}>0$ moves the AMM’s reserve ratio away from that of the other AMMs, which we call a divergent swap. This occurs when the initial vector of reserves satisfies $\frac{y_{i}}{x_{i}}\leq\frac{y_{-i}}{x_{-i}}$. In this case, asset $X$ is relatively less scarce locally than globally, causing the nGMM to pay a higher price $\Delta y^{nGMM}_{i}$ for $X$ than the CPMM price $\Delta y^{CPMM}_{i}$. Consequently, the GMM price $\Delta y^{GMM}_{i}$ coincides with the CPMM price. This is illustrated in Figure 3.
@@ 5.1 Arbitrage
There are no striclty profitable arbitrage opportunities between a set of AMMs that use the GMM.
@@
Furthermore, if traders strictly benefit from the switch, their terms of trade must improve, whereas if AMMs strictly benefit, their liquidity pools should grow, reducing slippage and improving terms of trade for future traders. In either case, better terms of trade should lead to increased trading and thus higher potential gains from the switch.
To assess the distribution of gains, we consider a stylized benchmark. First, we assume that all strictly profitable arbitrage opportunities are exploited by arbitrageurs and that traders always submit their trades to the AMM offering the most favorable terms of trade, a strategy we call optimal routing. As we shall see, this assumption aligns with empirical evidence.
Second, when all AMMs use the CPMM algorithm, the individual liquidity ratios after maximum arbitrage, $\frac{y_{i}}{x_{i}}$, must be equal across AMMs. However, this common ratio may depend on how arbitrageurs extract their profit. For instance, if arbitrageurs structure their swaps to realize profits exclusively in $Y$, ensuring that the net position of asset $X$ across all arbitrage swaps remains zero, the global liquidity ratio $\frac{\sum_{i}y_{i}}{\sum_{i}x_{i}}$ will decrease, leading to a corresponding decline in the common individual liquidity ratios $\frac{y_{i}}{x_{i}}$.
To abstract from this complication, we assume that arbitrage profit extraction preserves the global liquidity ratio, ensuring that the resulting individual liquidity ratios remain equal to the original global liquidity ratio. We refer to this assumption as balanced arbitrage.
@@
For cases where (10) applies and thus worse terms of trade may impact trading activity, there is a natural solution, which we describe next.
@@ 5.2 The GMM with Rebalancing
First, an important aspect of the GMM that we have not discussed but is now relevant is how swaps between AMMs using the GMM algorithm are priced. Up to this point, we have implicitly assumed that these trades do not occur. While such orders do not require special consideration under the CPMM, they exhibit a particular property in the nGMM: they do not affect the global aggregate $(x,y)$. Thus, for any swap $\Delta x$ of $X$ sent to an AMM, any corresponding $\Delta y$ is compatible with maintaining a constant aggregate product.
@@
Next, we introduce a modification of the GMM designed to guarantee traders an improvement over the CPMM.
Definition: The GMM algorithm with rebalancing is a global algorithm that applies the GMM algorithm to the resulting updated vector of reserves from executing the following algorithm iteratively until one of the conditions fails (assume wlog $\Delta x>0)$.
@@
(b) provides better terms of trade for the trader.
@@ 5.3 MEV Sandwich Attacks
By definition, the profit of ane elementary MEV sandwich is equal to $\widehat{\Delta^{F}x}_{s}-\widehat{\Delta x},$ for algorithm $s\in\{CPMM,GMM\}$, where:
@@
Thus, the min implies that the profits are weakly less under the GMM than under the CPMM. To show that they are strictly less we only need to show that in one of the mins, the second term is strictly less then the first. For that, just note that if it were not the first one, the following chaing of implications implies that it has to be the third one:
@@ 5.4 Impermanent Losses
We adopt a simple setting that we refer as the ideal benchmark. In the ideal benchmark, we assume that there are two AMMs only $1$ and $2$, with respectively reserves $(x_{1},y_{1})$ and $(x_{2},y_{2})$. As in the case of the CPMM, we also assume that the initial ratio of reserves of both AMMs is equal to the initial price $r_{0}$, i.e. $\frac{y_{1}}{x_{1}}=\frac{y_{2}}{x_{2}}=r$ and we model a change in market sentiment assuming an insider trader that submit trades to the AMM to maximize the value of its net final position at a the new price $r^{\prime}$.
@@
Finally, we assume without loss of generality that $\alpha\equiv\frac{x_{2}}{x_{1}+x_{2}}\leq 0.5$, which under the above assumptions means that AMM 1 is the AMM with largest initial market value $rx_{i}+y_{i}$.
@@ 6 Empirical Evaluation
We proceed with evaluating the theoretical results with simulations based on real trades. We consider transactions in two AMMs in the Ethereum network, Uniswap V2 and Sushiswap. [^8] We chose Uniswap V2 instead of V3 due to the similar fee structure that it has compared to Sushiswap for ease of exposition. Moreover, we assume that all fees return to liquidity pools for both AMMs and ignore the 0.25% that is reserved in the treasury for Sushiswap The datasets are used to illustrate each of the metrics discussed in previous sections: arbitrage profits, profist of MEV sandwich attacks and impermament losses. We flag arbitrage and sandwich operations using the identification of the ZeroMev database. [^9] https://zeromev.org/
We start by providing an assessment of the benefits of replacing the CPMM with the GMM in terms of the cost of arbitrage. We consider 3 periods for one of the most traded pairs USDT/WETH, two periods with high price variation and one with low. Moreover, we consider three other less traded pairs: ALCX/WETH, BLOCKS/WETH and VSP/WETH for periods that they display large arbitrage profits. We were not systematic on the section of pairs or periods beyond what we already explained. Table 1 presents the description of the samples and their characteristics. Table 6 confirms that routing is not far away from the optimal routing assumption that we introduced in Subsection 5.1. Except for Blocks and VSP, the cost of suboptimally routing orders only means a cost in the range of a few basis points.
Pairs WETH/ Ethereum Blocks Timeframe Transactions Description
USDT (1) 11,565,020 - 11,640,849 Aug 2021-12 days 114,974 tx High volatility
USDT (2) 12,390,294 - 12,467,913 May 2021-12 days 153,944 tx High volatility
USDT (3) 13,317,339 - 13,406,729 Oct 2021-12 days 59,156 tx Low volatility
ALCX 11,948,971 - 12,050,004 Jul 2021-20 days 21,481 tx Low Volatility / Unequal pools
BLOCKS 14,356,780 - 14,497,019 Mar 2022-20 days 4,297 tx High Volatility / Equal Pools
VSP 13,528,195 - 13,917,471 Nov 2021-20 days 2,681 tx High Volatility / Equal Pools
Table 1: Samples
Pairs WETH/ Traders losses compared to optimal routing / Volume of Trade
USDT (1) 0.03 %
USDT (2) 0.07%
USDT (3) 0.016 %
ALCX 0.15%
BLOCKS 1.7 %
VSP 2.3%
To estimate the potential gains of replacing the CPMM with GMM with respect to the cost of arbitrage we use the arbitrage profits as discussed in Subsection 5.1. We provide this measure of the cost of arbitrage with respect to the revenue fees of the AMMs which may be interpreted as the cost of liquidity provision by the AMMs. Our results show that the gains for a traded pair are modest ranging from 0.23% to 5.82%. For the other less traded pairs, the gains are significant which suggests that the GMM can be particularly effective for less traded pairs.
Pairs WETH/ Fees Arbitrage Arbitrage/Fees Ratio
USDT (1) 2,659 WETH 154.8 WETH 0.0582
USDT (2) 4,187 WETH 138.19 WETH 0.0330
USDT (3) 719.3 WETH 1.67 WETH 0.0023
ALCX 466.25 WETH 192.25 WETH 0.4124
BLOCKS 27.4 WETH 190.93 WETH 6.9701
VSP 21.18 WETH) 53.7 WETH 2.5368
@@
Table 2: Arbitrage Profits and Fees
Lastly, we provide an estimate on the size of arbitrage based on a larger sample of three months (blocks 18251371 to 18908876, 01/10/23 to 31/12/23) for Uniswap V2. For this period, a total of 71.6 million USD was collected in fees while 3.4 million was extracted from arbitrageurs. The problem seems to be much more prevalent in smaller/less traded pairs. Specifically, out of 1,833 pairs that experienced at least one arbitrage operation, $20\%$ exhibited Arbitrage Profits to Fees ratios greater than $10\%$. These pairs represented approximately $7.45\%$ of the total AMMs fees collected in the period, an approximate meausure of the relative volume traded. Specifically these AMMs collected 5.33 million USD in fees while they experienced 2.56 million USD of arbitrage profits. This confirms the observation that the GMM may be particularly useful for less traded pairs that tend to have large arbitrage costs, measured by arbitrage profits, relative to liquidity provision costs, measured by the AMMs’ revenue from fees.
Next we proceed with assessing assess the effectiveness of the GMM algorithm in reducing the profitability of sandwich operations. To do so we analyze all sandwich operations that occurred for the most traded pairs with WETH for the period 2021 and 2023. For the simulations, we use the reserves at the beginning of each block where a sandwich operation appears and simulate the entire block of transactions affecting the AMMs of interest, Uniswap v2 and Sushiswap. Then, we compare the profit of the sandwich operation under CPMM and GMM. For the counterpart AMM (Sushiswap), when simulating the GMM, we use its reserves at the start of the same block but keep them constant, disregarding trades that may have occurred in parallel with the sandwich operation. To the extend that the excluded trades might increase the reserves of the AMMs, and that larger reserves make sandwich operations less profitable, our measure of sandwich profits for the GMM is an upper bound. Some sandwich attacks identified by zeromev seem to be substantially more complex than the elementary MEV sandwichs we describe in Section 5. If this is the case, we measure the profits as in Section 5 replacing the frontrun operation operation in the data with the corresponding backrun operation that makes the attack an elementary MEV sandwich. [^10] Measuring the profits of general MEV sandwich positions is an open question in the literature as it is usually the case in reality that the MEV sandwich attack is combined with other operations to economize gas usage. We find a more general estimate of MEV profist is beyond the scope of this paper. If an operation after this adjustment becomes unprofitable for the CPMM, we discard it. Our simulations suggest that this exclusion does not introduce significant bias, as our simulated MEV operators’ profits explain approximately $80\%$ of the profits reported by zeromev. Our results, presented in Table 3, show that in all cases, the GMM design significantly reduces the profitability of sandwich operations and, in many cases, renders them unprofitable, which suggest a substantial reduction of MEV attacks.
Pairs WETH/ Profitable Sandwich Attacks (CPMM) Profitable Sandwich Attacks (GMM) Sandwich Profits (CPMM) Sandwich Profits (GMM)
USDT (Uni) 446 177 \$988,254 \$260,000
USDC (Uni) 240 90 \$853,872 \$140,000
DAI (Uni) 443 94 \$922,269 \$84,311
WBTC (Uni) 123 18 4 WETH / 20 BTC 0.1 WETH / 0.78 BTC
LINK (Uni) 464 200 8,949 LINK / 85 WETH 1,479 LINK / 9.86 WETH
LDO (Sushi) 1,152 1,114 139,336 LDO / 260 WETH 132,183 LDO / 245 WETH
CRV (Sushi) 659 473 140,345 CRV / 197 WETH 21,102 CRV / 90 WETH
Table 3: Simulations for Sandwich operations
Finally, similarly to arbitrage we present our findings on the size of sandwich for the same longer period considered for arbitrage operations. Again, we measure the cost of sandwich attacks for traders with the profitability for the attackers, and compare it to the cost of liquidity provision measured by the revenue of AMMs from fees. In that period and for all the AMMs that exhibited at least one sandwich operation, the cost of sandwich attacks for traders is 90 million USD, which is close to the cost of liquidity provision of 94.7 million USD. Furthermore, out of 11,188 pairs that exhibited a sandwich operation, 94% of them had a ratio of trader’s sandwich cost to liquidity cost of 10% or more. Those AMMs collected a total of 36 million USD in fees while sandwitch operators profited 89.8 million USD.
Finally, we conclude our analysis by computing the impermament losses using real data and compare the CPMM with a simulated counterfactual corresponding to the GMM. We use data from three pairs. The first pair USDT/WETH is a well traded pair where both Sushiswap and Uniswap have relatively deep and equal pools. The second pair ALCX/WETH was chosen to demonstrate how the GMM performs when the reserves of the two AMMs are highly concentrated (Sushiswap pocesses more than $99\%$ of the total reserves for the period). Lastly, VSP/WETH represents a pair with relatively shallow but equal pools.
We add a simulated trend to the original data. We consider two cases corresponding to a price increase of $50\%$ ($r=1.5$) and $100\%$ ($r=2$) for the whole period. To make the prices fit this trend, we take the real transactions and increase orders for WETH and decrease orders for the counterpart currency for each consecutive day to fit the intended price increase for the whole period. Specifically, for the first day we sample the average of demand for WETH, we increase it by 2% and draw new trades, for the second day we increase the average by 4% and so on. Table 4 describes the results.
Pairs WETH/ Reserve Distribution Initial Price of WETH Time (Block #) Trend I.L. Difference / I.L. Total
USDT 48 - 52% 3,484 Aug 2021 (Block: 13,317,339) r=1.5 3.5%
USDT 48 - 52% 3,484 Aug 2021 (Block: 13,317,339) r=2 5.2%
ALCX 0.01 - 99.9% 1.83 Jul 2021 (Block: 11,948,971) r=1.5 0.2%
ALCX 0.01 - 99.9% 1.83 Jul 2021 (Block: 11,948,971) r=2 0.3%
VSP 62 - 38% 656 Sep 2021 (Block: 13,528,471) r=1.5 4.3%
VSP 62 - 38% 656 Sep 2021 (Block: 13,528,471) r=2 6.5%
Table 4: Impermanent Loss Comparison

### v2

@@ 4 The Analysis of Global Market Makers
for $x_{i}\equiv x_{i}^{0}+\Delta x_{i}$ and $y_{i}\equiv y_{i}^{0}-\Delta y_{i}$. Now suppose a distribution of reserves $(\vec{x},\vec{y})$ such that the other AMMs have the same ratio of reserves as $i$, i.e. $\frac{y_{i}}{x_{i}}=\frac{y_{j}}{x_{j}}$, and that $i$ has the largest product of reserves, i.e. $x_{i}y_{i}=\max_{j}x_{j}y_{j}$. The first property means that no arbitrage opportunities are exploited before the trader trades and the second that AMM $i$ is the most convenient AMM for the trader. Thus, trade preserving means that the terms of trade for the trader cannot be worse than the CPMM, which together with the definition of CPMM means that after any trade the product of reserves of AMM $i$ cannot increase. Thus if the trader submits $\Delta x$, AMM $i$ ends up with $x^{0}_{i}$ units of $X$ and a number of units of $Y$ strictly lower than $y_{0}$, which contradicts that the algorithm is not exploitable. ∎
Consequently, we propose the following algorithm.
@@
To understand how the GMM determines terms of trade, it is useful to distinguish whether a swap $\Delta x_{i}>0$ moves the AMM’s reserve ratio away from that of the other AMMs, which we call a divergent swap. This occurs when the initial vector of reserves satisfies $\frac{y_{i}}{x_{i}}\leq\frac{y_{-i}}{x_{-i}}$. In this case, asset $X$ is relatively less scarce locally than globally, causing the nGMM to pay a higher price $\Delta y^{nGMM}_{i}$ for $X$ than the CPMM price $\Delta y^{CPMM}_{i}$. Consequently, the GMM price $\Delta y^{GMM}_{i}$ coincides with the CPMM price. This is illustrated in Figure 3.
@@
That the GMM is natural and not exploitable is direct. We check whether it is trade preserving once we study its properties and in particular whether it displays arbitrage opportunities.
@@ 5.1 Arbitrage
There are no strictly profitable arbitrage opportunities between a set of AMMs that use the GMM.
@@
The question that remains is whether the distribution of these profits is sufficient to make the GMM trade preserving. [^7] An additional benefit that we do not investigate here is that the fact that the product of reserves of the AMMs using GMM does not decrease and strictly increases with convergent orders should mean that their liquidity pools grow and thus reduce slippage and improve terms of trade for future traders. To do this exercise we first note a peculiar feature of arbitrage. Exploiting all arbitrage opportunities when all AMMs use the CPMM mean that the liquidity ratios must be equal across AMMs. The common ratio, however, may vary depending on how arbitrageurs extract their profit. For instance, if arbitrageurs structure their swaps to realize profits exclusively in $Y$, ensuring that the net position of asset $X$ across all arbitrage swaps remains zero, the global liquidity ratio $\frac{\sum_{i}y_{i}}{\sum_{i}x_{i}}$ will decrease, leading to a corresponding decline in the common individual liquidity ratios $\frac{y_{i}}{x_{i}}$. To avoid contemplating all possible ways of resolving arbitrage opportunities, we focus on the most elementary case in which arbitrage profit extraction preserves the global liquidity ratio, ensuring that the resulting individual liquidity ratios remain equal to the original global liquidity ratio. We refer to this assumption as balanced arbitrage.
@@
Still Proposition 8 seems to say that the GMM is not trade preserving. We are now going to consider this question more in detail. Note that the comparison in the definition of trade preserving requires that all arbitrage opportunities are exploited, and this is not contemplated in scenario (b) of Proposition 8. Proposition 3 seems to suggest that this is irrelevant, but it is not difficult to show that the GMM allows for zero profit arbitrage opportunities, and thus a marginal decrease of the GMM prices in the convergent zone will make them strictly profitable. It may be shows that this is sufficient to make this marginally modified GMM trade preserving. However, we believe that there is more interesting solution. Keep the GMM algorithm as it is and include in the GMM an additional algorithm that implements the zero profit arbitrage opportunities required to make the GMM trade preserving. This is what we do in the next section.
@@ 5.2 The GMM with Rebalancing
In this section, we shall show how to modify the GMM algorithm to implement automatically all the zero profit arbitrage trades necessary (and no more than them) to make it trade preserving.
Clearly, this require defining how the terms of trade of swaps between AMMs using the GMM algorithm are priced. All our results hold true if these swaps are implemented by an external intermediary, but in practical terms it may be more convenient to do so directly. Note that this is a feature that does not require special consideration for the CPMM prices. However the nGMM prices exhibit a particular property: any swap between AMMs do not affect the global aggregate $(x,y)$. Thus, for any swap $\Delta x$ of $X$ sent from one AMM to another one, any corresponding $\Delta y$ is compatible with maintaining a constant aggregate product.
@@
Next, we introduce the modification of the GMM that guarantees that trade preservation.
Definition: The GMM algorithm with rebalancing is an algorithm that applies the GMM algorithm to the resulting updated vector of reserves from executing the following algorithm iteratively until one of the conditions fails (assume wlog $\Delta x>0)$.
@@
(b) provides better terms of trade for the trader than (a).
@@
Corollary 2.
The GMM with rebalancing is trade preserving and not exploitable.
@@ 5.3 MEV Sandwich Attacks
By definition, the profit of an elementary MEV sandwich is equal to $\widehat{\Delta^{F}x}_{s}-\widehat{\Delta x}$,
for algorithm $s\in\{CPMM,GMM\}$, where:
@@
Thus, the min implies that the profits are weakly less under the GMM than under the CPMM. To show that they are strictly less we only need to show that in one of the mins, the second term is strictly less then the first. For that, just note that if it were not the first one, the following chain of implications means that it has to be the third one:
@@ 5.4 Impermanent Losses
We adopt a simple setting that we refer as the ideal benchmark. In the ideal benchmark, we assume that there are two AMMs only $1$ and $2$, with respectively reserves $(x_{1},y_{1})$ and $(x_{2},y_{2})$. As in the case of the CPMM, we also assume that the initial ratio of reserves of both AMMs is equal to the initial price $r^{\prime}$, i.e. $\frac{y_{1}}{x_{1}}=\frac{y_{2}}{x_{2}}=r^{\prime}$ and we model a change in market sentiment assuming an insider trader that submit trades to the AMM to maximize the value of its net final position at a the new price $r$.
@@
Finally, we assume without loss of generality that $\alpha\equiv\frac{x_{2}}{x_{1}+x_{2}}\leq 0.5$, which under the above assumptions means that AMM 1 is the AMM with largest initial market value $r^{\prime}x_{i}+y_{i}$.
@@ 6 Empirical Evaluation
We proceed to evaluate the relevance of our theoretical results using real data and simulations based on actual trades. We focus on transactions from 2023 in the most popular AMM on the Ethereum network, Uniswap V2, and uses the most elementary version of the CPMM algorithm. [^11] In 2023, Uniswap V2 facilitated trades in 165,566 different token pairs, with a total trading volume of \$88 billion. In comparison, Ethereum’s total trading volume was \$513 billion, while the combined trading volume across all blockchains was \$885 billion. Thus, Uniswap V2 accounted for 17% of Ethereum’s volume, and Ethereum processed 58% of the total volume across all blockchains. Source: Dune Analytics https://dune.com., [^12] Uniswap has produced more sophisticated version of the CPMM algorith called Uniswap v3 and Uniswap v4. They still use local reserves to determine the terms of trade. To construct our dataset, we extract Uniswap transactions from TheGraph, https://thegraph.com, and flag arbitrage and sandwich attacks on Ethereum using the identification provided by Zeromev, https://zeromev.org/, both ordered by block. We also obtain block-level WETH prices from Alchemy, https://www.alchemy.com. To express transaction values in U.S. dollars, we use the dollar prices of WETH and major stablecoins (DAI, USDC, and Tether), which allows us to convert approximately 90% of the transaction volume. Pairs without reliable price data are excluded from the analysis.
Table 1 summarises the importance of arbitrage and sandwich attacks in our database relative to the fees charged by AMMs to traders. Since these fees are essentially the compensation to liquidity providers for locking their tokens in the AMM, they can be interpreted as the cost of liquidity provision. The table thus provides a comparison of the costs of arbitrage and sandwich attacks relative to the cost of liquidity provision. The table shows that arbitrage is not particularly significant, but sandwich attacks are.
Metric Value
Fees \$261m
Arbitrage profits \$10m
Sandwich profits \$440m
Arbitrage profits over fees 3.8%
Sandwich profits over fees 168.2%
Crypto pairs s.t. % of trade volume
$\tfrac{\mbox{Arbitrage profits}}{\mbox{Total fees}}>50\%$ 0.8%
$\tfrac{\mbox{Sandwich profits}}{\mbox{Total fees}}>50\%$ 41.8%
@@
Table 1: Fees, arbitrage profits, and sandwich profits for all pairs traded on Ethereum in 2023.
Next, we use our dataset to simulate the profitability of the sandwich attacks recorded in 2023 under the hypothetical scenario in which Uniswap V2, instead of using the CPMM algorithm, employs the GMM algorithm. To this end, we focus on recorded sandwich attacks that are relatively simple to track, specifically those in which the frontrun trade is equal to the backrun trade. This includes elementary sandwich attacks but also more general cases, as we allow for several intermediate trades between the frontrun and backrun trades. This restriction reduces the observed sandwich profits from \$440 million to \$339 million.
One challenge in this exercise is tracking the reserves of the other AMMs that may be trading the same token pairs. Rather than relying on real-world reserve data, we run our algorithm under different hypothetical scenarios. In particular, we consider two broad cases. In the first, we assume that the reserves of the other AMMs are equal to $\beta$ times the reserves of Uniswap v2, and we consider different values of $\beta$. In this case, the attacker’s profits can be computed as:
$\left(\frac{(1+\frac{\delta+\hat{\delta}}{1+\beta})(1+\delta+\hat{\delta})}{\left(1+\hat{\delta}\right)(1+\delta+\hat{\delta})-\frac{\delta}{1+\beta}}-1\right)\hat{\Delta}x,$ (23)
where $x_{i}$ denotes the reserves in Uniswap of the asset submitted by the attacker, $\hat{\Delta}x$ is the amount of this asset submitted by the attacker, $\Delta x$ is the total amount submitted by all traders, and $\delta\equiv\frac{\Delta x}{x_{i}}$ and $\hat{\delta}\equiv\frac{\hat{\Delta}x}{x_{i}}$.
The results are presented in Table 2. We also include a column showing the percentage of cases in which the attacker’s profits decrease so significantly that they become negative. In the same table, we report the sandwich profits recorded by Zeromev and the simulated attacker profits under the CPMM algorithm (which corresponds to $\beta=0$). The table shows that our CPMM simulations produce results consistent with the data from Zeromev. It also shows that the attacker’s profits can be substantially reduced by switching from the CPMM to the GMM algorithm.
Algorithm Attackers’ profits % of profits<<0
Zeromev \$338.70m 4.1 %
CPMM \$333.70m 3.3 %
GMM ($\beta=0.01$) \$255.72m 6.7 %
GMM ($\beta=0.05$) \$188.65m 6.7 %
GMM ($\beta=0.1$) \$164.40m 6.8 %
GMM ($\beta=0.5$) \$121.96m 7.4 %
GMM ($\beta=1$) \$107.22m 7.9 %
GMM ($\beta=1.5$) \$99.63m 8.2 %
GMM ($\beta=10$) \$78.78m 9.5 %
Table 2: Sandwiches reported by Zeromev and simulations for 2023: attacker’s profits and percentage of cases with negative profits for different values of the fraction $\beta$ of Uniswap V2’s reserves held by other AMMs.
In the second case, we conduct a different exercise. We consider a counterfactual scenario in which the Uniswap V2 pool of reserves is evenly divided among $n$ AMMs, all using the GMM algorithm, and where each AMM computes global reserves by aggregating the reserves of all $n$ AMMs. We compute the attacker’s profits assuming that the attack and the trades had been directed to one of these hypothetical AMMs. A straightforward computation shows that these profits are given by:
$\left(\frac{(1+(\delta+\hat{\delta}))(1+n\delta+n\hat{\delta})}{(1+n\hat{\delta})(1+n\delta+n\hat{\delta})-\delta}-1\right)\hat{\Delta}x.$ (24)
The purpose of this exercise is to construct a counterfactual that allows us to evaluate the GMM without assuming any potential increase in system-wide liquidity.
Again, Table 3 shows a significant reduction in the profitability of sandwich attacks and a high fraction of cases in which these profits disappear.
Algorithm Attackers’ profits Attacks with profits<<0
Zeromev \$338.70m 4.1 %
CPMM \$333.70m 3.30%
GMM (n=2) \$44.26m 78.22%
GMM (n=3) \$27.09m 84.01%
GMM (n=4) \$19.58m 87.16%
GMM (n=5) \$15.37m 89.20%
Table 3: Sandwiches reported by Zeromev and simulations for 2023: attacker’s profits and percentage of attacks with negative profits for different values of the number $n$ of equal-sized AMMs into which the Uniswap V2 is split.
Finally, we conclude our analysis by using our dataset to compare the CPMM and the GMM in terms of impermanent losses. This is a complex task, as it requires isolating the effects of various events that influence AMM reserves over time, primarily liquidity injections, withdrawals, and fee revenue. To simplify the computation while still obtaining a realistic estimate, we rely on Equations (4) and (18). These formulae allow us to calculate impermanent losses as a function of the change in the relative price of the crypto assets in the AMM’s liquidity pool between two points in time, and the share of the AMM’s reserves relative to the global reserves across all AMMs considered in the GMM. The main caveat is that Equation (18) assumes only two AMMs are included in the GMM formula, and that a single trade moves the AMM’s reserves to the point where its marginal price matches the new relative price. Allowing for multiple intermediate trades can only increase the value of the final portfolio of the AMM, thereby reducing impermanent losses. Similarly, incorporating more AMMs into the formula can only improve the performance of the GMM. In this sense, our assumptions provide a conservative estimate of the potential gains from adopting the GMM in place of the CPMM.
Equation (7) requires the initial and final relative prices of the crypto assets as input. For each pair, we compute these prices by identifying, in our sequence of transactions, the first and last time the pair was traded in Uniswap V2.131313In this process, we drop pairs Pairs without reliable price data in US\$, pairs with less than two transactions and those with values so close to zero that rounding errors suggest dropping them from the analysis. These reduces the number of pairs from 165,566 to 124,898. The total impermanent losses computed for the CPMM algorithm using Equation (4) amount to \$615.3 million. Equation (18) also requires the parameter $\alpha$, which denotes the share of Uniswap’s reserves relative to the global reserves considered under the GMM. As previously discussed, accurately tracking the reserves of other AMMs trading the same token pairs is a complex task. For simplicity, we select illustrative values of $\alpha$. [^14] Recall that Equation (18) requires $\alpha\leq 0.5$. This corresponds to a counterfactual in which Uniswap V2 uses the GMM algorithm and incorporates the reserves of another AMM whose size is $\frac{1-\alpha}{\alpha}$ times that of Uniswap V2.
Finally, we distinguish between pairs with low and high price volatility. The later when the relative price from the beginning to the end of 2023 varies more than 10 times.
$\alpha$ Low Volatility High Volatility
$IL_{CPMM}-IL_{GMM}$ % of $IL_{CPMM}$ $IL_{CPMM}-IL_{GMM}$ % of $IL_{CPMM}$
$0.01$ \$12.5m 98.4% \$136.6m 22.7%
$0.05$ \$12.0m 94.5% \$93.8m 15.6%
$0.1$ \$11.3m 89.0% \$76.9m 12.8%
$0.25$ \$9.3m 73.2% \$53.8m 8.9%
$0.5$ \$6.1m 48.0% \$33.3m 5.5%
Table 4: Reduction in impermanent losses for different values of $\alpha$ under the GMM. Percentages are relative to the CPMM impermanent losses of \$12.7m (low volatility) and \$602.6m (high volatility) under the CPMM.
Table 4 shows that the reduction in impermanent losses is substantial and largely driven by volatile pairs. These cases likely correspond to pairs in which one of the crypto assets failed, causing its relative price to collapse to near zero. As a result, liquidity providers who had deposited funds in these pools lost almost all their capital. To shed light on the magnitude of these price shifts, Table 5 reports the fraction of token pairs that experienced price changes by factors of 10, 100, and 1000.
Low Volatility Pairs High Volatility Pairs
$\lambda=10$ 87.5 % 12.5 %
$\lambda=100$ 92.1 % 7.9 %
$\lambda=1000$ 93.1 % 6.9 %
Table 5: Percentage of pairs with price variations greater than a factor of $\lambda$, for different values of $T$. The total number of pairs is 124,898.
7 Implementation of the GMM
In this section, we overview the technical feasibility of our proposed algorithm, the Generalized Market Maker (GMM), and provide some statistics on the cost of using the algorithm in terms of gas consumption on the Ethereum blockchain.
The implementation of the GMM in Solidity, the coding lenguage of Ethereum, is publicly available in https://github.com/JonathanJimenezMunoz/GMM/tree/main [^15] We thank Jonathan Jimenez Muñoz for developing this implementation. The design is based on the architecture of Uniswap V2, the most widely used implementation of CPMM, and modifies only the components necessary to allow the algorithm to account for global reserves. Thus, the implementation inherits the robust and battle-tested foundation of Uniswap V2. Besides, the modular architecture of Uniswap V2 facilitates maintainability, integration with existing infrastructure and future updates.
The computational cost of the GMM, measured in gas consumption (the unit of computational effort on the Ethereum network) together with its cost in dollars, is reported in Table 6. We compare the original CPMM implementation with the modified GMM version for different numbers of external AMMs included in the global reserve calculation. As expected, gas usage increases with the number of external references, since each additional AMM requires an extra data retrieval and computation step. However, the increase in gas cost is moderate and consistent with the additional informational complexity introduced by the GMM. These results suggest that the GMM can be deployed in practice with manageable increases in gas consumption.
Algorithm Gas Consumption Cost
CPMM 102,438 \$1.61
GMM (1 other AMM) 115,676 \$1.82
GMM (2 other AMMs) 120,619 \$1.90
GMM (3 other AMMs) 125,091 \$1.97
GMM (4 other AMMs) 129,492 \$2.04
GMM (5 other AMMs) 133,916 \$2.11
GMM (6 other AMMs) 138,317 \$2.18
Table 6: Gas consumption and cost for CPMM and GMM with varying numbers of other AMMs. Prices in dollars using a gas cost of 5.764 Gwei and ETH price of \$2,726.29, as of May 27th, 2025.